# Goodness-of-fit test for multi-layer stochastic block models

### v1

@@ Abstract
Community detection in multi-layer networks is a fundamental task in complex network analysis across various areas like social, biological, and computer sciences. However, most existing algorithms assume that the number of communities is known in advance, which is usually impractical for real-world multi-layer networks. To address this limitation, we develop a novel goodness-of-fit test for the popular multi-layer stochastic block model. The test statistic is derived from a normalized aggregation of layer-wise adjacency matrices. Under the null hypothesis that a candidate community count is correct, we establish the asymptotic normality of the test statistic using recent advances in random matrix theory. This theoretical foundation enables a computationally efficient sequential testing algorithm to determine the number of communities. Numerical experiments on simulated and real-world multi-layer networks demonstrate the accuracy and efficiency of our approach in estimating the number of communities.
@@ 1 Introduction
In single-layer networks modeled by the SBM model, the problem of determining the number of communities has been addressed through various statistical methods. To estimate $K$, notable approaches include network cross-validation [6, 29], Bayesian inference or composite likelihood Bayesian information criterion [34, 43, 16], likelihood ratio tests [46, 32], spectrum of the Bethe Hessian matrices [24, 18], and goodness-of-fit tests [3, 25, 9, 17, 22, 48]. For brief reviews, see [22]. However, multi-layer networks introduce unique challenges due to their layered complexity and intra-layer heterogeneity. Consequently, single-layer estimation techniques can not be immediately applied to estimate the number of communities for multi-layer networks. Therefore, specialized methodologies are required to handle the aggregation of information across multiple layers while preserving statistical tractability. To address this, this paper aims to propose an efficient method for estimating $K$ in multi-layer networks under the multi-layer SBM model. Our contributions are threefold:
  • 1.
    A sequential hypothesis testing algorithm: We develop a novel spectral-based goodness-of-fit test that leverages a normalized aggregation of layer-wise adjacency matrices. In detail, we introduce a test statistic based on the proposed normalized aggregation matrix to distinguish between adequate community structures (null hypothesis $H_{0}:K=K_{0}$) and underfitting (alternative hypothesis $H_{1}:K>K_{0}$) without requiring prior knowledge of $K$. The normalization ensures robustness to variations in edge probabilities across layers, a critical feature for real-world networks where layer-specific connectivity patterns can differ significantly.
  • 2.
@@
    Theoretical guarantees: Under mild regularity conditions, we establish the asymptotic normality of the test statistic under the null hypothesis. This result is derived using recent advances in random matrix theory, particularly the analysis of linear spectral statistics for generalized Wigner matrices. Notably, our theoretical analysis does not depend on the specific community detection algorithm used to estimate the community assignments. Instead, it only requires that the estimated communities satisfy certain misclustering error bounds—a condition that is achievable by many existing methods (e.g., bias-adjusted spectral clustering developed in [27]). This generality ensures the applicability of our framework across diverse algorithmic choices. To our knowledge, this is the first work to establish rigorous theoretical guarantees for estimating the number of communities under the multi-layer SBM.
  • 3.
@@
    Empirical validation: Extensive experiments on synthetic and real-world multi-layer networks demonstrate the high accuracy of our approach in recovering the number of communities.
@@
We organize the rest of this paper as follows. In Section 2, we formalize the multi-layer stochastic block model and outline the sequential hypothesis testing framework for determining the number of communities. Section 3 introduces the ideal normalized aggregation matrix and derives the asymptotic normality of the ideal test statistic under the null hypothesis. Section 4 proposes a feasible test statistic and establishes its asymptotic properties. The sequential testing algorithm is presented in Section 5. Sections 6 and 7 validate the method through extensive numerical experiments on synthetic and real-world multi-layer networks, respectively. We conclude with discussions and future extensions in Section 8. Technical proofs of theoretical results are provided in the Appendix.
@@ 3 Ideal test statistic and its asymptotic normality
There exists $\delta>0$ such that $B_{\ell,kl}\in[\delta,1-\delta]$ for all $\ell,k,l$.
@@
4 Proposed test statistic and its asymptotic normality
@@
5 Hypothesis testing algorithm
@@
In this section, we present our normalized aggregation spectral test (NAST) algorithm, a sequential goodness-of-fit procedure that determines the number of communities in a multi-layer stochastic block model without prior knowledge of $K$. Starting with $K_{0}=1$, our NAST repeatedly invokes a community detection routine to obtain an estimated community label vector, estimates layer-wise connectivity and probability matrices, forms the normalized aggregation matrix, and evaluates the cubic-trace test statistic $T$. The loop terminates when the test statistic falls inside the acceptance region of the standard normal distribution, at which point the current $K_{0}$ is returned as the estimated number of communities. In this paper, we use the bias-adjusted SoS algorithm of [27] as the community detection technique, because its mis-clustering error satisfies $m=o(\sqrt{n})$ under mild conditions, thereby guaranteeing that Assumption 3 holds and the asymptotic null distribution $N(0,1)$ is preserved. The details of our NAST method is summarized in Algorithm 2, where $z_{1-\alpha/2}$ denotes the $(1-\alpha/2)$-quantile of the standard normal distribution $N(0,1)$. In this paper, we adopt the conventional significance level $\alpha=0.05$, so $z_{1-0.05/2}=z_{0.975}\approx 1.96$.
@@
1:Adjacency matrices $\{A_{\ell}\}_{\ell=1}^{L}$, significance level $\alpha$
@@
5:  Obtain $\{\hat{P}_{\ell}\}^{L}_{\ell}$ via Algorithm 1.
6:  Compute $T$ via Equation (5).
7:  if $|T|<z_{1-\alpha/2}$ then
@@ 6 Numerical experiments
This section validates the theoretical properties of the proposed test statistic and evaluates the performance of the proposed NAST algorithm. We design experiments to verify the asymptotic normality of the test statistic $T$ under $H_{0}$, examine the size and power of $T$, and assess NAST’s accuracy in estimating the true number of communities $K$ across diverse multi-layer networks. All simulations use the bias-adjusted SoS algorithm [27] for community detection, with significance level $\alpha=0.05$ in NAST. Multi-layer network generation follows the multi-layer SBM framework in Definition 1, and the community assignments are generated by letting each node belong to each community with equal probability for all experiments.
@@
Experiment 3: Accuracy of NAST in estimating $K$. We assess NASA’s accuracy for true $K\in\{1,2,3,4,5\}$ under various conditions. Networks ($n=1000$, $L=10$) use connectivity matrices $B_{\ell,kl}=\rho(0.3+\epsilon_{\ell}+0.4\cdot\mathbf{1}(k=l))$, where $\epsilon_{\ell}\sim\text{Uniform}(-0.1,0.1)$ introduces distinct perturbations per layer and block, with $\rho\in\{0.01,0.05,0.1,0.15,0.2,\\ 0.25,0.3\}$ controlling network’s sparsity. We increase $K_{0}$ sequentially. Table 2 shows the proportion of correct estimates $\hat{K}=K$ over 200 trials. Based on the numerical results, we see that the proposed NAST algorithm demonstrates robust performance in estimating the true number of communities $K$ across varying sparsity levels $\rho$, particularly for $\rho\geq 0.05$. At the lowest sparsity level ($\rho=0.01$), estimation accuracy is highly sensitive to $K$: while accuracy remains high for $K=1$ (93.0%) and $K=2$ (92.0%), it drops drastically to 1.0% for $K=3,4$ and 0.5% for $K=5$, indicating that extreme sparsity impedes reliable community recovery in more complex structures. However, as $\rho$ increases moderately to $0.05$, accuracy sharply improves across all $K$, exceeding 92.0% in all cases.
Table 2: Proportion of correct estimates of $K$ over 200 trials at different sparsity levels $\rho$.
Sparsity level ($\rho$)
0.01 0.05 0.1 0.15 0.2 0.25 0.3
$K=1$ 0.930 0.920 0.950 0.945 0.955 0.975 0.945
$K=2$ 0.920 0.945 0.960 0.965 0.950 0.950 0.940
$K=3$ 0.010 0.950 0.970 0.945 0.920 0.940 0.960
$K=4$ 0.010 0.955 0.950 0.960 0.940 0.965 0.960
$K=5$ 0.005 0.940 0.990 0.970 0.935 0.940 0.965
Table 3: Average values of the goodness-of-fit test statistic $T$ over 200 Monte Carlo trials for different candidate community numbers $K_{0}$ when the true number of communities $K\in\{1,2,3,4,5\}$. The first $T$ value satisfying $|T|<z_{1-\alpha/2}\approx 1.96$ (i.e., the null hypothesis $H_{0}:K=K_{0}$ is accepted) for each true $K$ is highlighted in bold. The settings of network parameters $(n,L,\theta,\{B_{\ell}\}^{L}_{\ell=1})$ are the same as Experiment 3, with the sparsity parameter $\rho$ being $0.1$.

### v2

@@ Abstract
Community detection in multi-layer networks is a fundamental task in complex network analysis across various areas like social, biological, and computer sciences. However, most existing algorithms assume that the number of communities is known in advance, which is usually impractical for real-world multi-layer networks. To address this limitation, we develop a novel goodness-of-fit test for the popular multi-layer stochastic block model based on a normalized aggregation of layer-wise adjacency matrices. Under the null hypothesis that a candidate community count is correct, we establish the asymptotic normality of the test statistic using recent advances in random matrix theory; conversely, we prove its divergence when the model is underfitted. This dual theoretical foundations enable two computationally efficient sequential testing algorithms to consistently determine the number of communities without prior knowledge. Numerical experiments on simulated and real-world multi-layer networks demonstrate the accuracy and efficiency of our approaches in estimating the number of communities.
@@ 1 Introduction
In single-layer networks modeled by the SBM model, the problem of determining the number of communities has been addressed through various statistical methods. To estimate $K$, notable approaches include network cross-validation [24, 25], Bayesian inference or composite likelihood Bayesian information criterion [26, 27, 28], likelihood ratio tests [29, 30], spectrum of the Bethe Hessian matrices [31, 32], and goodness-of-fit tests [33, 34, 35, 36, 37, 38]. For brief reviews, see [37]. However, multi-layer networks introduce unique challenges due to their layered complexity and intra-layer heterogeneity. Consequently, single-layer estimation techniques can not be immediately applied to estimate the number of communities for multi-layer networks. Therefore, specialized methodologies are required to handle the aggregation of information across multiple layers while preserving statistical tractability. To address this, we propose a novel, principled framework for estimating the number of communities $K$ in multi-layer networks under the multi-layer SBM model. Our contributions are threefold:
@@
  • A novel test statistic and its theoretical foundation: We introduce a new goodness-of-fit test statistic based on a normalized aggregation of the layer-wise adjacency matrices. This construction is key to handling heterogeneous edge probabilities across layers. For this statistic, we establish its asymptotic normality under the null hypothesis $H_{0}:K=K_{0}$ and prove its divergence under the alternative hypothesis $H_{1}:K>K_{0}$. These theoretical results are derived using recent advances in random matrix theory, particularly the analysis of linear spectral statistics for generalized Wigner matrices. Notably, our theoretical analysis does not depend on the specific community detection algorithm used to estimate the community assignments. Instead, it only requires that the estimated communities satisfy certain misclustering error bounds—a condition that is achievable by many existing methods (e.g., bias-adjusted spectral clustering developed in [8]). This generality ensures the applicability of our framework across diverse algorithmic choices. These properties provide the essential theoretical backbone for model specification testing.
  • Efficient sequential testing algorithms: Building directly upon the above theoretical results, we design two computationally efficient sequential testing algorithms: the normalized aggregation spectral test (NAST) and the sequential ratio-based NAST (SR-NAST). We prove that both algorithms consistently estimate the true $K$. To our knowledge, this is the first work to establish rigorous theoretical guarantees for estimating the number of communities under the multi-layer SBM.
@@
  • Empirical validation: Extensive experiments on synthetic and real-world multi-layer networks demonstrate the high accuracy of our approaches in recovering the number of communities.
@@
The rest of this paper is organized as follows. In Section 2, we formalize the multi-layer stochastic block model. Section 3 introduces the ideal test statistic and establishes its asymptotic properties. Section 4 presents the feasible statistic and its theoretical foundation. The two consistent sequential testing algorithms, NAST and SR-NAST, are detailed in Section 5. Their empirical performance is validated in Sections 6 and 7 using synthetic and real-world data, respectively. We conclude with discussions in Section 8. Technical proofs are provided in the Appendix.
@@ 3 Ideal test statistic and its asymptotic normality
There exists $\delta\in(0,\frac{1}{2})$ such that $B_{\ell,kl}\in[\delta,1-\delta]$ for all $\ell,k,l$.
@@
4 Proposed test statistic and its theoretical properties
@@
The following conditions are needed for the test’s power against the multi-layer SBM model.
Assumption 4.
Define the average connection matrix $\bar{B}=\frac{1}{L}\sum_{\ell=1}^{L}B_{\ell}$. There exists a constant $\eta>0$ such that
$\min_{1\leq k\neq k^{\prime}\leq K}\;\max_{1\leq l\leq K}\;\bigl|\bar{B}_{kl}-\bar{B}_{k^{\prime}l}\bigr|\geq\eta.$
Assumption 4 ensures that the true communities remain distinguishable in the average connectivity matrix $\bar{B}$. This is a mild, natural condition: it only requires that after aggregating across layers, every pair of communities differs in how they connect to at least one other community by a non‑vanishing amount $\eta$. It does not demand separation in every individual layer, making it suitable for real multi‑layer data where layers can be heterogeneous. In fact, this condition is the multi‑layer analogue of the classical row‑separation condition considered in Theorem 3.3 of [34].
Assumption 5.
When $K_{0}<K$, we require
$\frac{nL}{K^{4}}\to\infty\mathrm{~and~}\frac{nL^{3}}{K^{12}(\log n)^{3}}\to\infty\quad\text{as }n\to\infty.$
Assumption 5 is mild in typical applications because the number of communities $K$ is usually much smaller than the number of nodes $n$. In the simplest case where $K$ is bounded ($K=O(1)$), both conditions reduce to $nL\to\infty$ and $nL^{3}/(\log n)^{3}\to\infty$, which hold naturally. Even if $K$ grows slowly with $n$, the two conditions remain easily satisfied as long as $K$ does not increase too fast relative to $n$ and $L$. Thus, Assumption 5 describes a realistic scaling regime that covers most practical multi‑layer network scenarios.
The following theorem provides a lower bound of $|T|$ under the alternative model $K_{0}<K$.
Theorem 2 (Asymptotic power guarantee).
Suppose that Assumptions 15 hold, and the true number of communities $K$ is greater than the candidate $K_{0}$ (i.e., $K_{0}<K$). Then, there exists a constant $C>0$ such that
$\lim_{n\to\infty}\mathbb{P}\left(|T(K_{0})|\geq C\cdot\frac{(nL)^{3/2}}{K^{6}}\right)=1.$
Theorem 2 establishes the asymptotic power of the proposed goodness-of-fit test against the underfitting case $K_{0}<K$. This theorem shows that the test statistic $|T(K_{0})|$ grows at least as fast as $(nL)^{3/2}/K^{6}$ in probability. This growth rate is derived from the spectral norm of a signal submatrix that arises because, when $K_{0}<K$, at least one estimated community must contain nodes from two distinct true communities. The separation condition in Assumption 4 ensures that the aggregated connectivity profiles of these true communities differ by a non‑vanishing amount $\eta$, which translates into a deterministic signal of order $\sqrt{nL}/K^{2}$ in the normalized aggregation matrix $\widetilde{A}^{\text{agg}}$. After controlling the noise part via Lemma 7 given in the appendix, the cubic trace $T(K_{0})$ inherits a magnitude of order $(nL)^{3/2}/K^{6}$.
Together with Theorem 1, Theorem 2 ensures a sharp phase transition: $|T(K_{0})|$ is $O_{P}(1)$ when $K_{0}=K$ but diverges at a polynomial rate when $K_{0}<K$. This behavior is exploited in our sequential algorithms to consistently estimate $K$. Thus, Theorem 2 is not merely a technical bound; it is the cornerstone for proving the consistency of our sequential estimation algorithms designed in the next section.
5 Hypothesis testing algorithms and their consistency in the estimation of $K$
The asymptotic normality of $T(K_{0})$ under $H_{0}$ (Theorem 1) and its divergence under $H_{1}$ (Theorem 2) provide a rigorous foundation for estimating the number of communities $K$ via sequential testing. In this section, we present two practical algorithms: the Normalized Aggregation Spectral Test (NAST) and the Sequential Ratio-based Normalized Aggregation Spectral Test (SR-NAST). Both methods are computationally efficient, requiring only a single community detection step per candidate $K_{0}$, and both are provably consistent.
5.1 Normalized Aggregation Spectral Test (NAST)
@@
NAST directly implements the sequential goodness-of-fit test. Starting with $K_{0}=1$, our NAST repeatedly invokes a community detection routine to obtain an estimated community label vector, estimates layer-wise connectivity and probability matrices, forms the normalized aggregation matrix, and evaluates the cubic-trace test statistic $T$. The loop terminates when $|T|$ falls below a prescribed threshold $t_{n}$ (e.g., $\log n$), at which point the current $K_{0}$ is returned as the estimated number of communities. In this paper, we use the bias-adjusted SoS algorithm of [8] as the community detection technique, because its mis-clustering error satisfies $m=o(\sqrt{n})$ under mild conditions, thereby guaranteeing that Assumption 3 holds and the asymptotic null distribution $N(0,1)$ is preserved. The details of our NAST method is summarized in Algorithm 2.
@@
1:Adjacency matrices $\{A_{\ell}\}_{\ell=1}^{L}$, threshold $t_{n}$
@@
5:  Obtain $\{\hat{P}_{\ell}\}^{L}_{\ell}$ via Algorithm 1 using candidate community number $K_{0}$.
6:  Compute $T(K_{0})$ via Equation (5).
7:  if $|T(K_{0})|<t_{n}$ then
@@
13:return $\hat{K}$
The threshold $t_{n}$ must be chosen to separate the null distribution from the alternative. The following theorem provides explicit conditions on $t_{n}$ that guarantee consistency.
Theorem 3 (Consistency of NAST).
Suppose that Assumptions 15 all hold. Define the NAST estimator $\hat{K}$ as in Algorithm 2, where the threshold $t_{n}$ satisfies:
  • $t_{n}\to\infty$,
  • $t_{n}=o\!\left((nL)^{3/2}/K^{6}\right)$.
Then,
$\lim_{n\to\infty}\mathbb{P}\bigl(\hat{K}=K\bigr)=1.$
Theorem 3 follows directly from Theorems 1 and 2. The conditions on $t_{n}$ ensure that $t_{n}$ grows slower than the divergence rate under $H_{1}$ but faster than the bounded fluctuations under $H_{0}$. Thus, as $n\to\infty$, $\mathbb{P}(|T(K_{0})|>t_{n})\to 1$ for all $K_{0}<K$, while $\mathbb{P}(|T(K)|\leq t_{n})\to 1$. Consequently, NAST rejects all $K_{0}<K$ and accepts exactly at $K_{0}=K$ with probability approaching one. A simple valid choice is $t_{n}=\log n$, which satisfies both conditions under Assumption 5 because $(nL)^{3/2}/K^{6}$ grows much faster than $\log n$. In practice, one might consider thresholds of the form $c\log n$ for some constant $c>0$, but the asymptotic theory remains valid for any such threshold sequence satisfying conditions in Theorem 3.
5.2 Sequential Ratio-Based Normalized Aggregation Spectral Test (SR-NAST)
SR-NAST replaces the absolute test statistic in NAST with a ratio statistics. This method leverages the distinct asymptotic behavior of this ratio: it remains stochastically bounded when the candidate number of communities is below the true value but diverges to infinity when the candidate number equals the true value. This sharp phase transition allows the algorithm to reliably identify the true number of communities using a simple, slowly diverging threshold (e.g., $\log n$).
This method examines the ratio statistic defined below:
$\eta_{K_{0}}=|T(K_{0}-1)|/|T(K_{0})|\qquad{K_{0}=2,3,\ldots,K}.$ (6)
The behavior of the ratio statistic $\eta_{K_{0}}$ is characterized by the following theorem, which justifies the stopping rule.
Theorem 4 (Behavior of the ratio statistic).
Under Assumptions 1-5:
  1. 1.
    For $2\leq K_{0}\leq K-1$ (underfitting),
    $\eta_{K_{0}}=O_{P}(1).$
  2. 2.
    When $K_{0}=K$ (true model),
    $\eta_{K}\xrightarrow{P}\infty.$
Theorem 4 captures the phase transition in $\eta_{K_{0}}$. For $K_{0}\leq K-1$, both $|T(K_{0}-1)|$ and $|T(K_{0})|$ diverge at the same rate $(nL)^{3/2}/K^{6}$ by Theorem 2 and its proof, so their ratio is stochastically bounded. At $K_{0}=K$, the numerator |T(K1)||T(K-1)| diverges by Theorem 2 while the denominator $|T(K)|$ is $O_{P}(1)$ by Theorem 1, causing $\eta_{K}$ to diverge. This sharp change makes $\eta_{K_{0}}$ a natural diagnostic for detecting $K$.
SR-NAST sequentially computes $\eta_{K_{0}}$ and stops when it exceeds a simple threshold $t_{\mathrm{ratio},n}$ (e.g., $t_{\mathrm{ratio},n}=\log n$). Algorithm 3 provides the complete procedure.
With the result of Theorem 4, the consistency of SR-NAST follows immediately.
Theorem 5 (Consistency of SR-NAST).
Under Assumptions 1-5, define the SR-NAST estimator $\hat{K}$ as in Algorithm 3, where the threshold $t_{\mathrm{ratio},n}$ satisfies:
  • $t_{\mathrm{ratio},n}\to\infty$,
  • $t_{\mathrm{ratio},n}=o\!\left((nL)^{3/2}/K^{6}\right)$.
Then,
$\lim_{n\to\infty}\mathbb{P}(\hat{K}=K)=1.$
Theorem 4 implies that at $K_{0}=K$, $\eta_{K}>t_{\mathrm{ratio},n}$ with probability tending to one; while for $K_{0}=2,\dots,K-1$, $\eta_{K_{0}}$ is bounded and hence $\eta_{K_{0}}>t_{\mathrm{ratio},n}$ with probability tending to zero. The initial step ($K_{0}=1$) is handled by the same threshold: if $K=1$, Theorem 1 gives $|T(1)|=O_{P}(1)$, so $|T(1)|\leq t_{\mathrm{ratio},n}$ with high probability; if $K>1$, Theorem 2 gives $|T(1)|$ divergent, so $|T(1)|>t_{\mathrm{ratio},n}$ with high probability, and the algorithm moves to $K_{0}=2$. Thus, SR-NAST iterates through $K_{0}=1,2,\dots$ and stops exactly at $K_{0}=K$ with probability approaching one. The choice $t_{\mathrm{ratio},n}=\log n$ is convenient and satisfies requirements in Theorem 5.
Algorithm 3 Sequential Ratio-based Normalized Aggregation Spectral Test (SR-NAST)
1:Multi-layer adjacency matrices $\{A_{\ell}\}_{\ell=1}^{L}$, threshold $t_{\mathrm{ratio},n}>0$
2:Estimated number of communities $\hat{K}$
3:Set $K_{0}\leftarrow 1$
4:Compute $T(1)$ via Equation (5)
5:if $|T(1)|\leq t_{\mathrm{ratio},n}$ then
6:  $\hat{K}\leftarrow 1$
7:  return $\hat{K}$
8:end if
9:Set $K_{0}\leftarrow 2$
10:repeat
11:  Compute $T(K_{0})$ via Equation (5) with candidate number $K_{0}$
12:  Compute $\eta_{K_{0}}$ via Equation (6)
13:  if $\eta_{K_{0}}>t_{\mathrm{ratio},n}$ then
14:   $\hat{K}\leftarrow K_{0}$
15:  else
16:   $K_{0}\leftarrow K_{0}+1$
17:  end if
18:until $\eta_{K_{0}}>t_{\mathrm{ratio},n}$
19:return $\hat{K}$
Remark 2.
SR-NAST is demonstrably more robust than NAST due to fundamental differences in their stopping rules. While both algorithms employ the same test statistic $T(K_{0})$—whose asymptotic normality under the correct model ($H_{0}:K=K_{0}$) and divergence under underfitting ($H_{1}:K>K_{0}$) form the basis for sequential testing—their decision criteria differ critically. NAST stops when the absolute value $|T(K_{0})|$ falls below a predefined threshold $t_{n}$. This requires identifying the point where a diverging sequence transitions to an $O_{P}(1)$ variable, a task highly sensitive to the precise asymptotic calibration of $t_{n}$. In contrast, SR-NAST leverages the ratio statistic $\eta_{K_{0}}$, which is $O_{P}(1)$ for all $K_{0}<K$ but diverges in probability ($\eta_{K}\stackrel{{\scriptstyle P}}{{\rightarrow}}\infty$) precisely at $K_{0}=K$. This creates a sharp, discontinuous signal—a jump from stochastic boundedness to explosion—which is inherently easier and more reliable to detect than the continuous attenuation of $|T(K_{0})|$. Consequently, the termination of SR-NAST is less sensitive to the specific choice of its threshold $t_{\text{ratio},n}$ compared to the delicate balance required for $t_{n}$ in NAST.
@@ 6 Numerical experiments
This section validates the theoretical properties of the proposed test statistic and evaluates the performance of the proposed algorithms. We design experiments to verify the asymptotic normality of the test statistic $T$ under $H_{0}$, examine the size and power of $T$, and assess the accuracy of NAST and SR-NAST in estimating the true number of communities $K$ across diverse multi-layer networks. All simulations use the bias-adjusted SoS algorithm [8] for community detection. Multi-layer network generation follows the multi-layer SBM framework in Definition 1, and the community assignments are generated by letting each node belong to each community with equal probability for all experiments.
@@
Experiment 3: Accuracy of NAST and SR-NAST in estimating $K$. We assess the accuracy of our proposed methods for true $K\in\{1,2,3,4,5\}$ under various conditions. Networks ($n=1000$, $L=10$) use connectivity matrices $B_{\ell,kl}=\rho(0.3+\epsilon_{\ell}+0.4\cdot\mathbf{1}(k=l))$, where $\epsilon_{\ell}\sim\text{Uniform}(-0.1,0.1)$ introduces distinct perturbations per layer and block, with $\rho\in\{0.01,0.05,0.1,0.15,0.2,0.25,0.3\}$ controlling network’s sparsity. We increase $K_{0}$ sequentially. Table 2 shows the proportion of correct estimates $\hat{K}=K$ over 200 trials. Based on the numerical results, we see that the proposed NAST and SR-NAST algorithms demonstrate robust performance in estimating the true number of communities $K$ across varying sparsity levels $\rho$, particularly for $\rho\geq 0.1$. At the lowest sparsity level ($\rho=0.01$), estimation accuracies are highly sensitive to $K$: while accuracies remain high for $K=1$, they drop drastically for $K\geq 2$, indicating that extreme sparsity impedes reliable community recovery in more complex structures. However, as $\rho$ increases moderately to $0.1$, accuracies sharply improve across all $K$, exceeding 92% in all cases.
Table 2: Proportion of correct estimates of $K$ over 200 trials at different sparsity levels $\rho$ for NAST and SR-NAST algorithms.
Algorithm Sparsity level ($\rho$)
$K$ 0.01 0.05 0.1 0.15 0.2 0.25 0.3
1 NAST 1 1 1 1 1 1 1
SR-NAST 1 1 1 1 1 1 1
2 NAST 0.705 1 1 1 1 1 1
SR-NAST 0.545 1 1 1 1 1 1
3 NAST 0 1 1 1 1 1 1
SR-NAST 0 1 1 1 1 1 1
4 NAST 0 0.970 1 1 1 1 1
SR-NAST 0 0.845 1 1 1 1 1
5 NAST 0 0.120 1 1 1 1 1
SR-NAST 0 0.5 0.920 1 1 1 1
Table 3: Average values of the test statistic $T$ over 200 Monte Carlo trials for different candidate community numbers $K_{0}$ when the true number of communities $K\in\{1,2,3,4,5\}$. The first $T$ value satisfying $|T|<t_{n}$ with $t_{n}=\log n\approx 6.9078$ when $n=1000$ (i.e., the null hypothesis $H_{0}:K=K_{0}$ is accepted) for each true $K$ is highlighted in bold. The settings of network parameters $(n,L,\theta,\{B_{\ell}\}^{L}_{\ell=1})$ are the same as Experiment 3, with the sparsity parameter $\rho$ being $0.1$.